OpenStudy (anonymous):

in a 20-acre field, 300 bushels of wheat are produced. At this rate, how many bushels will be produced in a 50-acre field

OpenStudy (anonymous):

lengendarysadist im ready for help please help me

OpenStudy (anonymous):

Well \[\frac{ 300 }{ 20 }=15\] so that means for every acre, there is 15 bushels produced. 15*50=750 so 750 bushels

OpenStudy (anonymous):

20/300 = 50/W

OpenStudy (anonymous):